OpenStudy (anonymous):

Calculate the integral:

OpenStudy (amistre64):

i hope its a good one

OpenStudy (anonymous):

\[\int\limits_{0}^{\infty}|9e ^{-(1/4)t}-6e ^{-(1/2)t}|dt\]

OpenStudy (anonymous):

It is, it is.... in another topic i asked about the ≤

OpenStudy (across):

first, break it down into the sum of two integrals.

OpenStudy (across):

(or difference, in this case)

OpenStudy (across):

then factor out the constants.

OpenStudy (amistre64):

focus on the generic: e^(-x) type

OpenStudy (across):

\[\int e^{ax}dx=\frac{1}{a}e^{ax}+c\]

OpenStudy (anonymous):

i know that but with these sign | |

OpenStudy (amistre64):

\[\frac{d}{dx}(e^{-kx})=-k\ e^{-kx}\] so all your missing to match is the "-k" part

OpenStudy (amistre64):

eww.. absolute value integrals

OpenStudy (amistre64):

it might help to know that |a| = a, and |-a| = a

OpenStudy (amistre64):

id have to read up on the abs value stuff again

OpenStudy (amistre64):



OpenStudy (amistre64):

\[\int_{a}^{b}|f(x)|dx=\int_{a}^{c}-(f(x))dx+\int_{c}^{b}f(x)dx\]

OpenStudy (amistre64):

and of course using the imporper integration techniques for this one might be useful as well

OpenStudy (anonymous):

but guys.. i have a question... do i have to take the abs value of my integral? or is that the last thing i have to do?

OpenStudy (anonymous):

\[[-36e ^{-(1/4)t}+12e ^{-(1/2)t}]\] or will it be [+36e ^{-(1/4)t}+12e ^{-(1/2)t}]

OpenStudy (amistre64):

\[9e^{−(1/4)t} − 6e^{−(1/2)t}\] take the first derivative to find any critical points: \[\frac{-9}{4}e^{−(1/4)t} + \frac{6}{2}e^{−(1/2)t}=0\] after some algebra which I messed up on but the wolf gives: we get: t = 8 ln(2) - 4ln(3) but better yet, graphing this thing tells us there is no cusp in the interval so we just regard this as the + side http://www.wolframalpha.com/input/?i=%7C9e%5E%28-t%2F4%29-6e%5E%28-t%2F2%29%7C

OpenStudy (amistre64):

in other words, ignore the | | since they are pointless across the interval [3,inf)

OpenStudy (amistre64):

where did I get 3 from? doh!! still, the cusp, or the bad part, is not in your interval so just integrate it without the | |

OpenStudy (anonymous):

but amisre64? what do you mean... there are 2 ways to calculate this, but what is the right one? Take the abs. first and then fill in the boundaries or fill in the boundaries first and then take the abs. I will get the answer 24 or 48 whats the right one?

OpenStudy (amistre64):

24 is the right answer an absolute value function tends to have what is called a cusp; that is the pointy part that has no derivative. if the cusp was in our interval we would integrate it in 2 sets; but since the cusp is not in our interval we can go ahead and ignore the | | notation

OpenStudy (anonymous):

if 24 is the right answer... i have to fill in the boundaries and then take the abs. value of it

OpenStudy (amistre64):

then id go that route :) since wolfram gives us 24 as an answer

OpenStudy (anonymous):

hmm... oke.. weird why is my teacher saying that i have to take the +36 instead of -36... then he just gave me the wrong answer

OpenStudy (anonymous):

but thank you for your explanation amistre64 :)